what volume (in ml)of 0.2M H2SO4 solution should be mixed with the 40ml of 0.1M NaOH SOLUTIONsuchthat the resulting solution has concentration of H2SO4  as (6/55)M
 

Grade:12th Pass

1 Answers

Arun
5 years ago

We know that resultant molarity is given by

Resultant Molarity = (M1V1+M2V2)/(V1+V2)

we have M1 = 0.1, V1 = 40ml = 0.04 litre

M2 = 0.2, V2 = ?

Now putting the values in above formula we get,

(0.1x0.04 + 0.2xV2)/0.04+V2 = 6/55

=> 0.004+0.2V2 = (0.24+6V2)/55

=> 0.22+11V2 = 0.24+6V2

=> 5V2 = 0.02

=> V2 = 0.004 litre

Hence, V2 = 0.004 litre or 4ml